OpenStudy (anonymous):

Need some help om just a few please.solve x^2 -7x< or = to 0

OpenStudy (anonymous):

x(x-7)<o x=o x-7=0 x=7

OpenStudy (anonymous):

x(x-7)=o x=o x-7=0 x=7